# The Jacobi Factoring Circuit: Quantum Factoring with Near-Linear Gates and Sublinear Space and Depth

### v1

@@ 2.3 Computational Number Theory
However, there is a faster divide-and-conquer algorithm which was conceived by Schönhage [Sch71] and [BS96, solution to exercise 5.52], and later expounded by Möller [Möl08]. This algorithm runs in $\widetilde{O}(n)$ gates [^6] While these algorithms are usually formulated in the Turing machine model, they can be readily transformed into circuits at the expense of multiplicative $\mathsf{polylog}(n)$ overheads [PF79]. (and hence at most that much space and depth).
@@
Finally, we remark that it is well-known that fully factoring an $n$-bit integer $N$ in the special case where $N$ is a prime power can be done in classical $\mathsf{poly}(n)$ time. This is because of two straightforward facts: (a) we can find integers $A,k>1$ such that $N=A^{k}$ if they exist by computing $N^{1/k}$ for all possible values of $k$ (and $k_{\mathrm{max}}<\log N$ since $N=A^{k}\geq 2^{k}$); and (b) we can efficiently test whether $A$ is prime [AKS04].
@@
Shor’s algorithm [Sho97] was the first to show that arbitrary $n$-bit integers could be factored using quantum circuits of size $\widetilde{O}(n^{2})$. However, Shor’s algorithm does not benefit if $N$ has a small prime divisor or small squarefree part, like the classical algorithms by [Len87, Mul24] do. The same holds of Regev’s [Reg23] improvement on Shor’s algorithm to use $\widetilde{O}(n^{3/2})$ gates.
To the best of our knowledge, the only polynomial-time [^8] An alternative approach to factoring with quantum computers is to use quantum subroutines (e.g. Grover search) inside classical factoring algorithms. The benefits this yields, which can include sublinear qubit count and gains from special structure in $N$, come at the expense of superpolynomial gate count and depth [BHLV17, BBM17, MBV20]. quantum factoring circuit that benefits from special structure in $N$ is the aforementioned Jacobi factoring circuit by [LPDS12], combined with the near-linear time algorithms for computing Jacobi symbols [Sch71, BS96, Möl08]. Putting these constructions together yields a circuit of only $\widetilde{O}(n)$ gates and space for finding a factor of $N$ when $N$ is not squarefree. Viewed in this context, one of our contributions is showing that we can further drive down the space and depth of this circuit when the squarefree part $B$ of $N=A^{2}B$ is much smaller than $N$.
@@ 2.4.2 Gauss Sums
Here, we state results that essentially imply that the Jacobi symbol is appropriately “pseudorandom” for the purposes of our algorithm and that by [LPDS12]. We follow the lecture notes by Conrad [Con].
@@
This allows us to prove the final result we need. For this, we first need a preliminary lemma:
@@
Let $m\in\mathbb{N}$ be squarefree with $r$ prime divisors. Then we have
$\sum_{j\in\mathbb{Z}_{m}:\gcd(j,m)=1}\exp\left(\frac{2\pi ij}{m}\right)=(-1)^{r}.$
@@
It is well-known that the LHS and RHS are both equal to the Möbius symbol $\mu(m)$ [HW75, Theorem 271]. ∎
We next state the specific form of the Gauss sum bound that we will need. We refer the reader to Section 1.1 for an overview of where these sums of phases come from, and reiterate the intuition here. Informally, if we want to recover $Q$ given $N=P^{2}Q$ as input, we will end up with a superposition of several periodic signals with period $Q$. The below lemma (with $m=Q$) examines the result of applying a QFT to this superposition, and tells us that these signals will essentially interfere like a randomly chosen collection of periodic signals.
Lemma 2.17.
Suppose $m$ is odd and squarefree and $k\in\mathbb{Z}_{m}$ is such that $\gcd(k,m)=1$. Then for any $\sigma\in\left\{-1,1\right\}$, we have:
$\left|\sum_{j\in\mathbb{Z}_{m}:\left(\frac{j}{m}\right)=\sigma}\exp\left(-\frac{2\pi ijk}{m}\right)\right|=\frac{\sqrt{m}}{2}+O(1).$
Proof.
First, note by a simple change of variables that
$\sum_{j\in\mathbb{Z}_{m}:\left(\frac{j}{m}\right)=\sigma}\exp\left(-\frac{2\pi ijk}{m}\right)=\sum_{j\in\mathbb{Z}_{m}:\left(\frac{j}{m}\right)=\sigma^{\prime}}\exp\left(\frac{2\pi ij}{m}\right),$
where $\sigma^{\prime}=\sigma\cdot\left(\frac{-k}{m}\right)\in\left\{-1,1\right\}$. Denote the RHS of this expression by $S_{\sigma^{\prime}}$, then note that on the one hand, by Lemma 2.16, we have $S_{1}+S_{-1}=(-1)^{r}$. On the other hand, we have:
$|S_{1}-S_{-1}|$ $=\left|\sum_{j\in\mathbb{Z}_{m}}\left(\frac{j}{m}\right)\exp\left(\frac{2\pi ij}{m}\right)\right|$
$=\sqrt{m},$
by Theorem 2.15. So for any $\sigma\in\left\{-1,1\right\}$, we have:
$|S_{\sigma}|$ $=\frac{1}{2}\left|(S_{1}+S_{-1})+(S_{\sigma}-S_{-\sigma})\right|$
$=\frac{\sqrt{m}}{2}+O(1),$
@@ 3 Factoring Squarefull Integers
using $S:=S(\ell,n)$ ancilla qubits with $G(\ell,n)$ gates and $D(\ell,n)$ depth, for any positive integer $x$ such that $x<2^{\ell}$.
@@
gates, $D(\ell,n)+O(\ell)$ depth, and $S(\ell,n)+\ell+2$ qubits, and any classical pre/post-processing is polynomial-time.
Before we prove the theorem, notice that plugging in the Jacobi symbol algorithm due to [Sch71, BS96, Möl08] (discussed in Section 2.3) and simply setting $B_{\mathrm{max}}=N$ immediately yields the following corollary: [^9] We note that the original paper by [LPDS12] does not appear to state a result using Schönhage’s near-linear size Jacobi algorithm [Sch71]; rather, they work with the better-known quadratic-time algorithms for computing the Jacobi symbol. Nevertheless, we credit [LPDS12] with this result since this result does essentially follow directly from their analysis (they need only set $\ell=\Theta(n)$ in Theorem 3.1).
@@
  • The uniform superposition over $[1,2^{\ell}]$ can be initialized in depth $O(1)$ and gates $O(\ell)$, using no ancilla qubits.
@@
    Compute the Jacobi symbol $\left(\frac{x}{N}\right)$ in superposition, to obtain the following state:
    $\frac{1}{2^{\ell/2}}\sum_{x=1}^{2^{\ell}}\ket{x}\Ket{\left(\frac{x}{N}\right)}\ket{0^{S}}.$
@@
    Measure the register containing $\left(\frac{x}{N}\right)$ to obtain a value $\sigma\in\left\{-1,0,1\right\}$. If $\sigma=0$, abort and terminate. Otherwise, we proceed as follows:
@@
    • For the classical post-processing, use the continued fraction expansion of $\frac{x^{*}}{2^{\ell}}$ (as in Shor [Sho97]; see [HW75, Chapter X] for details) to find positive integers $X_{1}$ and $X_{2}$ such that $X_{2}\leq B_{\mathrm{max}}$ and $\left|\frac{x^{*}}{2^{\ell}}-\frac{X_{1}}{X_{2}}\right|$ is minimal. Output $X_{2}$ and terminate. (We will show that with probability $\Omega(1)$, we will in fact have $X_{2}=B$.)
@@
It now remains to prove the correctness of Algorithm 3.1. To do this, we first prove a preliminary technical lemma, then turn our attention to proving the theorem. Let us briefly motivate the need for this technical lemma. Ideally we would like to be able to say that for any $x$ in our superposition, we have $\left(\frac{x}{N}\right)=\left(\frac{x}{A}\right)^{2}\left(\frac{x}{B}\right)=\left(\frac{x}{B}\right)$, and hence after measuring we end up with a superposition over values $x$ such that $\left(\frac{x}{B}\right)=\sigma$. The problem is that this is only true if $\left(\frac{x}{A}\right)\in\left\{-1,1\right\}$. This is true most of the time, but there will be a small fraction of inputs $x$ (specifically, those that share common factors with $A$ but not $B$) such that $\left(\frac{x}{B}\right)\in\left\{-1,1\right\}$ but $\left(\frac{x}{A}\right)=0\Rightarrow\left(\frac{x}{N}\right)=0$. The following lemma informally says that because there are not many such $x$’s, we can safely ignore this technicality: even though our algorithm prepares the state |$\ket{\psi_{2}}$, we can safely pretend that it in fact prepares the simpler state |$\ket{\psi_{1}}$ (by a trace distance argument).
@@
Let $M=\Omega(Bn^{c})$ be some parameter. Let $\sigma\in\left\{-1,1\right\}$, define |$\ket{\psi_{1}}$ and |$\ket{\psi_{2}}$ to be the following unnormalized states:
|$\ket{\psi_{1}}$ $=\sum_{\begin{subarray}{c}1\leq x\leq M\\ \left(\frac{x}{B}\right)=\sigma\end{subarray}}\ket{x}$
|$\ket{\psi_{2}}$ $=\sum_{\begin{subarray}{c}1\leq x\leq M\\ \left(\frac{x}{B}\right)=\sigma\\ \gcd(x,N)=1\end{subarray}}\ket{x}.$
Then the corresponding normalized states are $O(n^{(1-c)/4})$-close in trace distance. Moreover, we have |$\lVert\ket{\psi_{1}}\rVert_{2}^{2}=\frac{M\varphi(B)}{2B}(1+o(1)).$ (Note that these two states may not be identical, since there could be integers $x$ that share a common factor with $N$ but not $b$.)
@@
|$\lVert\ket{\psi_{1}}\rVert_{2}^{2}$ =|$=\left|\left\{x\in[1,M]:\left(\frac{x}{B}\right)=\sigma\right\}\right|$
$=\sum_{j\in[1,B-1]:\left(\frac{j}{B}\right)=\sigma}\left|\left\{x\in[1,M]:x\equiv j\bmod{B}\right\}\right|$
@@
$=\sum_{j\in[1,B-1]:\left(\frac{j}{B}\right)=\sigma}\left(\frac{M-j}{B}-\frac{1-j}{B}+O(1)\right)$
@@
$=\frac{M\varphi(B)}{2B}(1+o(1))\text{ (noting that $M=\omega(B)$)}.$
@@
|$\lVert\ket{\psi_{1}}-\ket{\psi_{2}}\rVert_{2}^{2}$ =|$=\left|\left\{x\in[1,M]:\left(\frac{x}{B}\right)=\sigma\text{ and}\gcd(x,N)>1\right\}\right|$
$\leq\sum_{i=1}^{r}\left|\left\{x\in[1,M]:\left(\frac{x}{B}\right)=\sigma\text{ and }p_{i}\mid x\right\}\right|$
$=\sum_{i=1}^{r}\sum_{j\in[1,B-1]:\left(\frac{j}{B}\right)=\sigma}\left|x\in[1,M]:x\equiv j\bmod{B}\text{ and }x\equiv 0\bmod{p_{i}}\right|.$
For each $i$, $B$ and $p_{i}$ are relatively prime so the conditions $x\equiv j\bmod{B}$ and $x\equiv 0\bmod{p_{i}}$ can be condensed into one condition that $x\equiv x_{i}\bmod{Bp_{i}}$ by the Chinese Remainder Theorem (for some unique residue $x_{i}\in\mathbb{Z}_{Bp_{i}}$). With this in mind, we continue as follows:
|$\lVert\ket{\psi_{1}}-\ket{\psi_{2}}\rVert_{2}^{2}$ $=\sum_{i=1}^{r}\sum_{j\in[1,B-1]:\left(\frac{j}{B}\right)=1}\left|x\in[1,M]:x\equiv j\bmod{B}\text{ and }x\equiv 0\bmod{p_{i}}\right|$
$=\sum_{i=1}^{r}\sum_{j\in[1,B-1]:\left(\frac{j}{B}\right)=1}\left|x\in[1,M]:x\equiv x_{i}\bmod{Bp_{i}}\right|$
$=\sum_{i=1}^{r}\sum_{j\in[1,B-1]:\left(\frac{j}{B}\right)=\sigma}\left(\left\lfloor\frac{M-x_{i}}{Bp_{i}}\right\rfloor-\left\lceil\frac{1-x_{i}}{Bp_{i}}\right\rceil+1\right)$
$\leq\sum_{i=1}^{r}\sum_{j\in[1,B-1]:\left(\frac{j}{B}\right)=\sigma}\left(\frac{M-1}{Bp_{i}}+1\right)$
[…]
$=O\left(\frac{M\varphi(B)}{Bn^{c-1}}\right)\text{ (since $M=\Omega(Bn^{c})$)}.$
@@
$O\left(\sqrt{\frac{\lVert\ket{\psi_{1}}-\ket{\psi_{2}}\rVert_{2}}{\lVert\ket{\psi_{1}}\rVert_{2}}}\right)$ $\leq O\left(\sqrt[4]{\frac{M\varphi(B)/(Bn^{c-1})}{M\varphi(B)/(2B)}}\right)$
$\leq O\left(n^{(1-c)/4}\right),$
@@
We now prove Theorem 3.1. Our proof breaks down into a few steps: we will first show that the measured Jacobi symbol $\sigma$ will be nonzero most of the time, before setting up some notation and writing out the state computed by the algorithm after the QFT. We then lower bound the amplitude this state places on certain values $y\in[0,2^{\ell}-1]$, and use this to complete the proof.
Step 1: disposing of the case where $\sigma=0$ and the algorithm aborts.
To do this, we will simply argue that this event occurs with probability $o(1)$:
Lemma 3.4.
We have $\Pr[\sigma=0]=o(1)$.
Proof.
Note that the probability of $\sigma=0$ is exactly
$\frac{1}{2^{\ell}}\left|\left\{x\in[1,2^{\ell}]:\gcd(x,N)>1\right\}\right|.$
Our analysis to bound this is similar to the proof of Lemma 3.3. Let $p_{1},\ldots,p_{r}$ be the distinct primes dividing $N$. We must have $r\leq n$ and $p_{i}\geq n^{c}$ for all $i$. With this in mind, we have:
$\frac{1}{2^{\ell}}\left|\left\{x\in[1,2^{\ell}]:\gcd(x,N)>1\right\}\right|$ $\leq\frac{1}{2^{\ell}}\sum_{i=1}^{r}\left|\left\{x\in[1,2^{\ell}]:x\equiv 0\bmod{p_{i}}\right\}\right|$
$=\frac{1}{2^{\ell}}\sum_{i=1}^{r}\left\lfloor\frac{2^{\ell}}{p_{i}}\right\rfloor$
$\leq\sum_{i=1}^{r}\frac{1}{p_{i}}$
$\leq\frac{r}{n^{c}}$
$\leq n^{1-c},$
which is $o(1)$ as claimed. ∎
Remark.
Note that aborting in the $\sigma=0$ case is unnecessary if the goal is just to output some nontrivial factor of $N$; if we get $\sigma=0$, the resulting superposition would be over integers $x$ such that $\gcd(x,N)>1$, so all we need to do is measure the $x$ register and take the $\gcd$ of the result with $N$ (the probability of the bad case that we end up with $N$ can be shown to be $o(1)$). For convenience, we rule out the possibility of $\sigma=0$ (which happens with small probability) and focus on establishing a stronger claim about what our algorithm accomplishes, namely that it either outputs $B$ or a prime dividing $N$. This in turn enables us to prove that this algorithm can be used for fully factoring a large class of integers in Section 5.
Step 2: notation and setup.
By Lemma 3.4, we may assume going forward that $\sigma\in\left\{-1,1\right\}$. In this case, we will use the observation that for any $x$, such that $\gcd(x,N)=1$, we have
$\left(\frac{x}{N}\right)=\left(\frac{x}{A}\right)^{2}\left(\frac{x}{B}\right)=\left(\frac{x}{B}\right),$
since $\left(\frac{x}{A}\right)\in\left\{-1,1\right\}$. Thus in this case, we have a state proportional to
|$\ket{\psi_{2}}=\sum_{\begin{subarray}{c}1\leq x\leq 2^{\ell}\\ \left(\frac{x}{B}\right)=\sigma\\ \gcd(x,N)=1\end{subarray}}\ket{x}.$
Noting that $2^{\ell}\geq B^{2}=\Omega(Bn^{c})$, we can use Lemma 3.3 to change this to the state
|$\ket{\psi_{1}}=\sqrt{\frac{(2+o(1))B}{2^{\ell}\varphi(B)}}\sum_{\begin{subarray}{c}1\leq x\leq 2^{\ell}\\ \left(\frac{x}{B}\right)=\sigma\end{subarray}}\ket{x},$
@@
$\sqrt{\frac{(2+o(1))B}{2^{2\ell}\varphi(B)}}\sum_{y=0}^{2^{\ell}-1}\left(\sum_{\begin{subarray}{c}1\leq x\leq 2^{\ell}\\ \left(\frac{x}{B}\right)=\sigma\end{subarray}}\exp\left(-\frac{2\pi ixy}{2^{\ell}}\right)\right)\ket{y}.$ (2)
@@
Step 3: lower bounding the amplitude on |$\ket{y}$.
@@
$\left|\sum_{\begin{subarray}{c}1\leq x\leq M\\ \left(\frac{x}{B}\right)=\sigma\end{subarray}}\exp\left(-\frac{2\pi ixy}{M}\right)\right|\geq\Omega(q\sqrt{B}).$

### v3

@@ 2.2 The Jacobi Symbol
At various points, we will use the notation $j_{b}(a)\in\left\{-1,1\right\}$ to denote a value that is guaranteed to be $\left(\frac{a}{b}\right)$ when $\gcd(a,b)=1$ and can be arbitrary otherwise (since as we will see below, in this case the Jacobi symbol would be 0).
@@ 2.3 Computational Number Theory
However, there is a faster divide-and-conquer algorithm which was conceived by Schönhage [Sch71] and [BS96, solution to exercise 5.52], and expounded Thull and Yap [TY90] and Möller [Möl08]. This algorithm runs in $\widetilde{O}(n)$ gates [^7] While these algorithms are usually formulated in the Turing machine model, they can be readily transformed into circuits at the expense of multiplicative $\mathsf{polylog}(n)$ overheads [PF79]. (and hence at most that much space and depth).
@@
Since we consider algorithms for factorizing integers of the form $p^{r}q$ in the less-studied regime where $q$ is small, we note that there has been extensive classical lattice-based cryptanalysis for factoring integers of this form [BDH99, CFRZ16] with other special constraints. Specifically, we can factor integers of the form $p^{r}q^{s}$ in polynomial time provided that $\max(r,s)\geq\mathsf{poly}(\log p)$. We do not believe these algorithms extend to our setting; roughly, these algorithms seem to be effective in factoring $N$ when $N$ has a prime factor $p$ such that (a) $N$ is divisible by a relatively large power of $p$; and (b) $p$ is still not too small relative to $N$. In our setting where $N=p^{2}q$ and $q$ is very small, neither of these is the case. Regardless, we emphasize that there has been little classical cryptanalysis for factoring integers of the specific form we consider (beyond Mulder’s aforementioned algorithm [Mul24]), and we leave this important direction to future work.
Finally, we remark that it is well-known that completely factoring an $n$-bit integer $N$ in the special case where $N$ is a prime power can be done in classical $\mathsf{poly}(n)$ time. This is because of two straightforward facts: (a) we can find integers $A,k>1$ such that $N=A^{k}$ if they exist by computing $N^{1/k}$ for all possible values of $k$ (and $k_{\mathrm{max}}<\log N$ since $N=A^{k}\geq 2^{k}$); and (b) we can efficiently test whether $A$ is prime [AKS04].
@@
Shor’s algorithm [Sho97] was the first to show that arbitrary $n$-bit integers could be factored using quantum circuits of size $\widetilde{O}(n^{2})$. However, Shor’s algorithm does not benefit if $N$ has a small prime divisor or small squarefree part, like the classical algorithms by [Len87, Mul24] do. The same holds for Regev’s [Reg25] improvement on Shor’s algorithm to use $\widetilde{O}(n^{3/2})$ gates. A work by Ekerå and Håstad [EH17], later built upon by Chevignard et al. [CFS24], achieved a constant factor improvement in circuit costs that holds specifically for RSA integers (the product of two primes of roughly the same size), but the asymptotic scaling of the algorithms was unchanged.
To the best of our knowledge, the only polynomial-time [^9] An alternative approach to factoring with quantum computers is to use quantum subroutines (e.g. Grover search) inside classical factoring algorithms. The benefits this yields, which can include sublinear qubit count and gains from special structure in $N$, come at the expense of superpolynomial gate count and depth [BHLV17, BBM17, MBV20]. quantum factoring circuit that benefits by more than constant factors from special structure in $N$ is the aforementioned Jacobi factoring circuit by [LPDS12], combined with the near-linear time algorithms for computing Jacobi symbols [Sch71, TY90, BS96, Möl08]. Putting these constructions together yields a circuit of only $\widetilde{O}(n)$ gates and space for finding a factor of $N$ when $N$ is not squarefree. Viewed in this context, one of our contributions is showing that we can further drive down the space and depth of this circuit when the squarefree part $B$ of $N=A^{2}B$ is much smaller than $N$.
@@ 2.4.2 Gauss Sums
Here, we state results that essentially imply that the Jacobi symbol is appropriately “pseudorandom” for the purposes of our algorithm and that of [LPDS12]. We follow the lecture notes by Conrad [Con].
@@
This allows us to prove the specific form of the Gauss sum bound that we will need. We refer the reader to Section 1.1 for an overview of where these sums of phases come from, and reiterate the intuition here. Informally, if we want to recover $Q$ given $N=P^{2}Q$ as input, we will end up with a superposition of several periodic signals with period $Q$. The below lemma (with $m=Q$) examines the result of applying a QFT to this superposition, and tells us that these signals will essentially interfere like a randomly chosen collection of periodic signals.
@@
Suppose $m$ is odd and squarefree and consider any $k\in\mathbb{Z}_{m}$. Then we have:
$\left|\sum_{j\in\mathbb{Z}_{m}}\left(\frac{j}{m}\right)\exp\left(-\frac{2\pi ijk}{m}\right)\right|=\begin{cases}\sqrt{m},\text{ if $\gcd(k,m)=1$,}\\ 0,\text{ else.}\end{cases}.$
@@
First, we address the case where $\gcd(k,m)=1$. Let $\chi$ denote the Jacobi symbol mod $m$. In this case we make a simple change of variables:
$\sum_{j\in\mathbb{Z}_{m}}\left(\frac{j}{m}\right)\exp\left(-\frac{2\pi ijk}{m}\right)$ $=\sum_{j\in\mathbb{Z}_{m}}\left(\frac{-j/k}{m}\right)\exp\left(\frac{2\pi ij}{m}\right)$
$=\left(\frac{-k}{m}\right)\cdot G(\chi),$
and now the conclusion is immediate from Theorem 2.15 and Lemma 2.13.
It remains to address the case where $\gcd(k,m)=d>1$. Let $k^{\prime}=k/d$ and $m^{\prime}=m/d$. If $k\equiv 0\pmod{m}$, the conclusion is immediate from Corollary 2.5, so we may assume $m^{\prime}>1$. Since $m$ is squarefree, we have $\gcd(d,m^{\prime})=1$ and thus we can use the Chinese Remainder Theorem to associate a value $j\in\mathbb{Z}_{m}$ with its residues $j\bmod{m^{\prime}}$ and $j\bmod{d}$. Bearing this in mind, we have:
$\sum_{j\in\mathbb{Z}_{m}}\left(\frac{j}{m}\right)\exp\left(-\frac{2\pi ijk}{m}\right)$ $=\sum_{j\in\mathbb{Z}_{m}}\left(\frac{j}{m^{\prime}}\right)\left(\frac{j}{d}\right)\exp\left(-\frac{2\pi ijk^{\prime}}{m^{\prime}}\right)$
$=\sum_{j_{1}\in\mathbb{Z}_{d}}\sum_{j_{2}\in\mathbb{Z}_{m^{\prime}}}\left(\frac{j_{2}}{m^{\prime}}\right)\left(\frac{j_{1}}{d}\right)\exp\left(-\frac{2\pi ij_{2}k^{\prime}}{m^{\prime}}\right)$
$=\left(\sum_{j_{1}\in\mathbb{Z}_{d}}\left(\frac{j_{1}}{d}\right)\right)\cdot\left(\sum_{j_{2}\in\mathbb{Z}_{m^{\prime}}}\left(\frac{j_{2}}{m^{\prime}}\right)\exp\left(-\frac{2\pi ij_{2}k^{\prime}}{m^{\prime}}\right)\right)$
[…]
$=0,$
@@ 3 Factoring Squarefull Integers
using $S:=S(\ell,n)$ ancilla qubits with $G(\ell,n)$ gates and $D(\ell,n)$ depth, for any positive integer $x$ such that $x<2^{\ell}$. As in Definition 2.3, $j_{N}(x)\in\left\{-1,1\right\}$ is such that $j_{N}(x)=\left(\frac{x}{N}\right)$ when $\gcd(x,N)=1$ and can be arbitrary for other $x$.
@@
gates, $D(\ell,n)+O(\ell)$ depth, and $S(\ell,n)+\ell$ qubits, and any classical pre/post-processing is polynomial-time.
Before we prove the theorem, notice that plugging in the Jacobi symbol algorithm due to [Sch71, TY90, BS96, Möl08] (discussed in Section 2.3) and simply setting $B_{\mathrm{max}}=N$ immediately yields the following corollary: [^10] We note that the original paper by [LPDS12] does not appear to state a result using Schönhage’s near-linear size Jacobi algorithm [Sch71]; rather, they work with the better-known quadratic-time algorithms for computing the Jacobi symbol. Nevertheless, we credit [LPDS12] with this result since this result does essentially follow directly from their analysis (they need only set $\ell=\Theta(n)$ in Theorem 3.1).
@@
  • The uniform superposition over $[1,2^{\ell}]$ can be initialized in depth $1$ and gates $\ell$, using no ancilla qubits.
@@
    Compute the function $j_{N}(x)$ in superposition, to obtain the following state:
    $\frac{1}{2^{\ell/2}}\sum_{x=1}^{2^{\ell}}j_{N}(x)\ket{x}\ket{0^{S}}.$
@@
  6. 6.
    Finally, for the classical post-processing, use the continued fraction expansion of $\frac{x^{*}}{2^{\ell}}$ (as in Shor [Sho97]; see [HW75, Chapter X] for details) to find positive integers $X_{1}$ and $X_{2}$ such that $X_{2}\leq B_{\mathrm{max}}$ and $\left|\frac{x^{*}}{2^{\ell}}-\frac{X_{1}}{X_{2}}\right|$ is minimal. Output $X_{2}$ and terminate. (We will show that with probability $\Omega(1)$, we will in fact have $X_{2}=B$.)
@@
It now remains to prove the correctness of Algorithm 3.1. To do this, we first prove a preliminary technical lemma, then turn our attention to proving the theorem. The need for this technical lemma is twofold:
  • We would like to argue that we can safely ignore inputs $x$ in the superposition where $\gcd(x,N)>1$, so we will have $j_{N}(x)\neq 0=\left(\frac{x}{N}\right)$.
  • Even when $\gcd(x,N)=1$, we would ideally be able to say that for any $x$ in our superposition, we have $\left(\frac{x}{N}\right)=\left(\frac{x}{A}\right)^{2}\left(\frac{x}{B}\right)=\left(\frac{x}{B}\right)$, and hence after measuring we end up with a superposition over values $x$ of $\left(\frac{x}{B}\right)\ket{x}$. The problem is that this is only true if $\left(\frac{x}{A}\right)\in\left\{-1,1\right\}$. This is true most of the time, but there will be a small fraction of inputs $x$ (specifically, those that share common factors with $A$ but not $B$) such that $\left(\frac{x}{B}\right)\in\left\{-1,1\right\}$ but $\left(\frac{x}{A}\right)=0\Rightarrow\left(\frac{x}{N}\right)=0$.
The following lemma informally says that because there are not many $x$’s where either of the above issues come up, we can safely ignore these technicalities: even though our algorithm prepares the state |$\ket{\psi_{2}}$, we can safely pretend that it in fact prepares the simpler state |$\ket{\psi_{1}}$ (by a trace distance argument).
@@
Define |$\ket{\psi_{1}}$ and |$\ket{\psi_{2}}$ to be the following unnormalized states:
|$\ket{\psi_{1}}$ $=\sum_{1\leq x\leq M}\left(\frac{x}{B}\right)\ket{x}$
|$\ket{\psi_{2}}$ $=\sum_{1\leq x\leq M}j_{N}(x)\ket{x}.$
Then the corresponding normalized states are $O(n^{(1-c)/4}\cdot(\log n)^{1/4})$-close in trace distance. Moreover, we have |$\lVert\ket{\psi_{1}}\rVert_{2}^{2}=\frac{M\varphi(B)}{B}(1+o(1)).$ (Note that these two states may not be identical, for the aforementioned reasons.)
@@
|$\lVert\ket{\psi_{1}}\rVert_{2}^{2}$ =|$=\left|\left\{x\in[1,M]:\gcd(x,B)=1\right\}\right|$
$=\sum_{j\in[1,B-1]:\gcd(x,B)=1}\left|\left\{x\in[1,M]:x\equiv j\bmod{B}\right\}\right|$
@@
$=\sum_{j\in[1,B-1]:\gcd(j,B)=1}\left(\frac{M-j}{B}-\frac{1-j}{B}+O(1)\right)$
@@
$=\frac{M\varphi(B)}{B}(1+o(1))\text{ (noting that $M=\omega(B)$)}.$
@@
|$\lVert\ket{\psi_{1}}-\ket{\psi_{2}}\rVert_{2}^{2}$ $\leq O(1)\cdot\left|\left\{x\in[1,M]:\gcd(x,N)>1\right\}\right|$
$\leq O(1)\cdot\sum_{i=1}^{r}\left|\left\{x\in[1,M]:p_{i}\mid x\right\}\right|$
$\leq O(1)\cdot\sum_{i=1}^{r}\frac{M}{p_{i}}$
$\leq O\left(\frac{M}{n^{c-1}}\right)\text{ ($r\leq n$)}.$
@@
$O\left(\sqrt{\frac{\lVert\ket{\psi_{1}}-\ket{\psi_{2}}\rVert_{2}}{\lVert\ket{\psi_{1}}\rVert_{2}}}\right)$ $\leq O\left(\sqrt[4]{\frac{M/n^{c-1}}{M\varphi(B)/B}}\right)$
$\leq O\left(n^{(1-c)/4}\cdot\left(\frac{B}{\varphi(B)}\right)^{1/4}\right)$
$\leq O\left(n^{(1-c)/4}\cdot(\log\log B)^{1/4}\right)$
$\leq O\left(n^{(1-c)/4}\cdot(\log n)^{1/4}\right),$
@@
We now prove Theorem 3.1. Our proof breaks down into a few steps: we will first set up some notation and write out the state computed by the algorithm after the QFT. We then lower bound the amplitude this state places on certain values $y\in[0,2^{\ell}-1]$, and use this to complete the proof.
Step 1: notation and setup.
After computing the function $j_{N}(x)$, we have the state
$\frac{1}{2^{\ell/2}}\sum_{x=1}^{2^{\ell}}j_{N}(x)\ket{x}\ket{0^{S}}.$
We can now use Lemma 3.3 to change this to the state
|$\ket{\psi_{1}}=\sqrt{\frac{(1+o(1))B}{2^{\ell}\varphi(B)}}\sum_{1\leq x\leq 2^{\ell}}\left(\frac{x}{B}\right)\ket{x},$
@@
$\sqrt{\frac{(1+o(1))B}{2^{2\ell}\varphi(B)}}\sum_{y=0}^{2^{\ell}-1}\left(\sum_{\begin{subarray}{c}1\leq x\leq 2^{\ell}\end{subarray}}\left(\frac{x}{B}\right)\exp\left(-\frac{2\pi ixy}{2^{\ell}}\right)\right)\ket{y}.$ (2)
@@
Step 2: lower bounding the amplitude on |$\ket{y}$.
@@
$\left|\sum_{1\leq x\leq M}\left(\frac{x}{B}\right)\exp\left(-\frac{2\pi ixy}{M}\right)\right|\geq\Omega(q\sqrt{B}).$